# FormulaReasoning: A Dataset for Formula-Based Numerical Reasoning

### v5

@@ 4 Experimental Setup
We implemented a wide range of baseline methods for handling the questions within FormulaReasoning, including prompting LLMs to use zero-shot CoT (Wei et al., 2022; Kojima et al., 2022)) and training a formula retriever to be incorporated into RAG solutions. In addition, we developed two methods to improve the reasoning capabilities of fine-tuned models with fewer than 7B parameters: one that divides the reasoning process into formula generation, parameter extraction, and numerical calculation, and the other that leverages data enhancement. We also used preference learning to improve the reasoning capabilities of small models.
@@ 4.1 Dataset Split
We divided FormulaReasoning into three subsets: training, HoF (Homologous Formulas) test, and HeF (Heterologous Formulas) test, comprising 3958, 410, and 383 questions, respectively. All formulas in the HoF test set appeared in the training set, while in the HeF test set, each question required at least one formula not seen in the training set. This division was designed to evaluate the generalizability of fine-tuned models on new formulas.
@@ 4.2 Evaluated Methods
Following Kojima et al. (2022), we incorporated “Let’s think step by step” (i.e. CoT) into a zero-shot prompt to guide LLMs in generating reasoning steps. The prompt is in Appendix A.2.1. We conducted experiments on DeepSeek-R1 (DeepSeek-AI et al., 2025), GPT-4o, Qwen3, GLM-4-flash (GLM et al., 2024), Qwen-max, as well as small models including Qwen2.5-14B/7B (Yang et al., 2024b) and Llama3.1-8B (Meta, 2024).
@@
We found that directly prompting models having fewer than 7B parameters did not produce satisfactory results, so we performed supervised fine-tuning of small models. We conducted experiments on Qwen2.5-Math-7B-Instruct and Qwen2.5-Math-1.5B-Instruct.
Unlike large models, small models did not exhibit proficient performance in numerical extraction and calculation. In order to augment the reasoning capabilities of small models, we further explored the following two improvements: CoT-SFT and DA.
  • CoT-Supervised Fine-Tuning (CoT-SFT) We decomposed the reasoning process into three steps. First, we instructed the model to generate relevant formulas. The parameter names within the formulas were then extracted, allowing the model to extract their numerical values and units from the question. Finally, the formulas and extracted numerical values and units of parameters were provided to the Numbat calculator to obtain an answer, thus relieving the model of numerical calculation.
  • Data Augmentation (DA) We augmented the training set with an LLM (Qwen-max). Firstly, we used a few-shot approach to prompt the LLM to generate new question-answer pairs. The LLM-generated explanations were then normalized. Finally, we verified the correctness of the normalized reasoning steps using Numbat. Details are in Appendix A.2.3.
@@ 4.3 Evaluation Metrics
We utilized Numbat to evaluate the prediction generated by each model against the gold-standard answer. A prediction was deemed correct if the relative error, (prediction -- gold) / gold, was less than 1%. We used accuracy, which is the proportion of questions correctly answered, as our metric.
@@ 5 Experimental Results
We present the evaluation results on the Chinese version of FormulaReasoning. The results on the English version are in Appendix A.3.1, which are consistent with those on the Chinese version.
@@
Model HoF test HeF test Avg. DeepSeek-R1 78.54 78.33 78.44 GPT-4o 65.12 61.36 63.24 Qwen3 63.17 61.62 62.40 GLM-4-flash 38.29 43.08 40.69 Qwen-max 74.63 72.85 73.74 Qwen2.5-14B 69.27 67.62 68.45 Qwen2.5-7B 44.63 48.04 46.34 Llama3.1-8B 11.22 10.44 10.83 Human 93.49 90.47 92.03
@@ 5.1 Results of LLMs with CoT Prompts
The evaluation results for LLMs with CoT prompts are given in Table 5. DeepSeek-R1 achieved the highest accuracy, outperforming its closest competitor by an average of 4.70 points. Among models with sizes below 20B, Qwen2.5-14B demonstrated commendable performance. However, none of the models aligned with human performance, an average over 108 students from a high school who each received a gift after completing 7–8 questions as their in-class exercise (approved by IRB), characterizing FormulaReasoning as an unsolved challenge for LLMs. We provide an error analysis in Appendix A.4.
PRM scores to assess the reasoning steps generated by LLMs are included in Appendix A.3.2. We calculated the Pearson correlation coefficient between accuracy and PRM score (one/multi-step). The high values of 0.90/0.77 indicated their strong positive correlation.
5.2 Results of LLMs with PoT Prompts
A comparison between CoT and PoT prompts is in Appendix A.3.3. PoT did not lead to better results. Therefore, the bottleneck for current LLMs on FormulaReasoning was mainly in their ability to apply formulas, rather than calculation, reflecting a unique challenge posed by our dataset.
Table 6: Accuracy of LLMs with CoT prompts and formula retriever.
Model HoF test DeepSeek-R1 78.73    + formula retriever 80.00 $+1.27$ Qwen2.5-14B 70.13    + formula retriever 72.66 $+2.53$
@@ 5.3 Results of LLMs with Formula Retriever
The results of two representative LLMs augmented with our formula retriever are shown in Table 6. We found that incorporating retrieved formulas into the prompts improved the performance of LLMs. For DeepSeek-R1 and Qwen2.5-14B, their accuracy increased by 1.27 and 2.53, respectively, on the HoF test set. Note that it would be meaningless to carry out this experiment on the HeF test set where the required formulas were not included when training the retriever.
Our investigation found that the top five formulas retrieved often included irrelevant ones, as the number of formulas required varied for different questions. The presence of extraneous formulas might introduce noise. Therefore, there was considerable room for further research in retrieving structured formulas from a database.
@@ 5.4 Results of Supervised Fine-Tuned Models
The results of supervised fine-tuned small models are included in Appendix A.3.4. Both CoT-SFT and DA helped improve their accuracy. In particular, Qwen2.5-Math-7B-Instruct with CoT-SFT outperformed GPT-4o and approximated Qwen-max, demonstrating the effectiveness of our fine-tuning method: delegating numerical calculations to a calculator and focusing on CoT reasoning.
@@ 5.5 Results of DPO (including Results on Other Numerical Reasoning Datasets)
As shown in Appendix A.3.5, DPO with the preference data derived from the training set of FormulaReasoning generally improved the accuracy of 1.5B–7B models on both FormulaReasoning and other numerical reasoning datasets: GSM8K, MATH, SVAMP, and GaoKao2023-en. These models exhibited much lower accuracy on FormulaReasoning than those datasets. The results reflected the challenging nature of our dataset and its additional value as a potential DPO resource.
6 Conclusion and Limitations
We introduced FormulaReasoning, a bilingual dataset for formula-based numerical reasoning. For each question, we annotated its complex reasoning steps with normalized physics formulas. We constructed a consolidated formula database after merging equivalent formulas, serving as an external knowledge base to be used in RAG. We used our dataset to evaluate LLMs of various sizes, RAG, fine-tuned small models, and DPO methods, demonstrating its challenging nature as an unsolved task. Our findings indicate substantial room for improvement of existing LLMs in formula-based numerical reasoning, or more generally, in multi-step reasoning guided by domain knowledge.
@@ 6 Conclusion and Limitations
Limitations One limitation of our work is that although FormulaReasoning provides step-level supervision information (i.e., formulas), due to the diversity of reasoning paths and formula expressions, we only used PRM to perform step-level process evaluation. Another limitation is that our dataset is focused on physics. We chose junior high school physics because it is not too hard to be understood by ordinary people, which benefited our annotation and evaluation efforts. It is possible to explore formula-based numerical reasoning in other domains such as chemistry and engineering.
@@ A.2.2 Formula Retriever
Each question might correspond to multiple correct formulas, and we ensured that the same question did not appear twice in the same batch when loading the data. Based on the implementation of Chinese-BERT-wwm-base, we tested the retrieval performance on the HoF test set and found that Recall@5 reached 97.69%.
@@ A.2.5 Implementation Details
We accessed DeepSeek-R1, [^4] https://api-docs.deepseek.com/ GPT-4o, Qwen3 (qwen3-235b-a22b version), GLM-4-flash, [^5] https://open.bigmodel.cn/ Qwen-max(qwen-max-2025-01-25 version), Qwen2.5-14B/7B [^6] https://bailian.console.aliyun.com/ and Llama3.1-8B through API calls with their default hyper-parameters. Qwen2.5-Math-7B/1.5B-Instruct were fine-tuned on NVIDIA GeForce RTX 3090 GPUs with Huggingface Transformers and Pytorch 2.3, where we used a learning rate of 1e-4 and a batch size of 8 to fine-tune them with LoRA (Hu et al., 2021) with r=8, alpha=32, and tested the models after training for 50 epochs. We set temperature=0 for direct answer generation and set the maximum output length to 1,024. In the case of CoT-SFT, which directly outputted formulas along with numerical values and units of parameters, we set temperature=1.0 and top_p=0.9, and if the output contained formatting errors, we allowed the small model to retry up to 3 times until a correctly formatted output was generated. Training Qwen2.5-Math-7B-Instruct and Qwen2.5-Math-1.5B-Instruct used 24 and 12 hours, respectively.
@@ A.3.1 English Version of FormulaReasoning
The evaluation results for LLMs with CoT prompts on the English FormulaReasoning dataset are presented in Table 7. DeepSeek-R1 achieved the highest average accuracy of 74.62, surpassing other models such as Qwen-max (69.70). Their ranking was generally consistent with that on the Chinese version presented in Section 5.1.
@@
Model HoF test HeF test Avg.
DeepSeek-R1 74.82 74.42 74.62
GPT-4o 66.03 60.61 63.32
Qwen3 68.54 69.45 69.00
GLM-4-flash 39.67 40.41 40.04
Qwen-max 69.83 69.57 69.70
Qwen2.5-14B 69.36 64.45 66.91
Qwen2.5-7B 32.07 36.57 34.32
Llama3.1-8B 12.83 11.00 11.92
A.3.2 PRM Scores
Tables 8 and 9 show the PRM scores in one-step and multi-step settings, respectively.
Table 8: PRM score (one-step) of LLMs with CoT prompts.
Model HoF test HeF test Avg.
DeepSeek-R1 0.6113 0.6318 0.6216
GPT-4o 0.6458 0.5424 0.5941
Qwen3 0.7034 0.7067 0.7051
GLM-4-flash 0.4127 0.3881 0.4004
Qwen-max 0.8656 0.8488 0.8572
Qwen2.5-14B 0.6766 0.6530 0.6648
Qwen2.5-7B 0.5692 0.5537 0.5615
Llama3.1-8B 0.1643 0.1397 0.1520
Table 9: PRM score (multi-step) of LLMs with CoT prompts.
Model HoF test HeF test Avg.
DeepSeek-R1 0.6777 0.6994 0.6886
GPT-4o 0.8405 0.8337 0.8371
Qwen3 0.8074 0.8090 0.8082
GLM-4-flash 0.7327 0.7291 0.7309
Qwen-max 0.9747 0.9719 0.9733
Qwen2.5-14B 0.8439 0.8186 0.8313
Qwen2.5-7B 0.8362 0.8311 0.8337
Llama3.1-8B 0.3582 0.3780 0.3681
A.3.3 PoT Prompts
@@
GPT-4o (CoT) 65.12 61.36 63.24
GPT-4o (PoT) 53.41 52.22 52.82

### v6

@@ 4 Experimental Setup
We implemented a wide range of reasoning approaches for handling the questions within FormulaReasoning, with a focus on understanding how different methodological choices impact performance on formula-based numerical reasoning tasks. Our investigation includes analyzing various prompting strategies for language models, developing specialized fine-tuning techniques, and incorporating external knowledge through retrieval-augmented methods. Specifically, we developed two complementary approaches to enhance reasoning capabilities: one that decomposed the reasoning process into formula generation, parameter extraction, and numerical calculation, and another that leveraged data augmentation techniques. We also explored preference learning to refine reasoning processes. This comprehensive experimental setup allows us to compare the strengths and limitations of different architectural and methodological choices for formula-based reasoning.
@@ 4.1 Dataset Split
We divided FormulaReasoning into three subsets: training, HoF (Homologous Formulas) test, and HeF (Heterologous Formulas) test, comprising 4,524, 413, and 387 questions, respectively. All formulas in the HoF test set appeared in the training set, while in the HeF test set, each question required at least one formula not seen in the training set. This division was to evaluate the generalizability of fine-tuned models on new formulas.
@@ 4.2 Evaluated Methods
Following Kojima et al. (2022), we incorporated “Let’s think step by step” (i.e. CoT) into a zero-shot prompt to guide LLMs in generating reasoning steps. The prompt is in Appendix A.2.1.
We conducted experiments across a diverse spectrum of models including cost friendly small-scale models including GLM-4.5-Air (GLM et al., 2024), ERNIE-4.5-21B-A3B Baidu-ERNIE-Team (2025), Gemma-3N-E4B Gemma (2025), Llama-3.3-8B-Instruct (Meta, 2024), Mistral-Small-3.2-24b-Instruct (Mistral, 2025), GPT-oss-20B (Agarwal et al., 2025), Gemini-2.5-Flash-Lite (Comanici et al., 2025) and Seed-OSS-36B-Instruct Team (2025), alongside high-cost LLMs such as GPT-4o (OpenAI, 2024), GLM-4-plus (GLM et al., 2024), Qwen3-235B-A22B-Thinking-2507 (Yang et al., 2025), DeepSeek-R1 (DeepSeek-AI et al., 2025), o3-mini (OpenAI, 2025b), and GPT-5 (OpenAI, 2025a). This comprehensive evaluation allowed us to assess capabilities across different model scales while identifying persistent challenges in formula reasoning.
@@
We found that directly prompting models having fewer than 8B parameters did not produce satisfactory results, so we performed supervised fine-tuning of small models. We conducted experiments on Qwen2.5-Math-7B-Instruct and Qwen2.5-Math-1.5B-Instruct.
Unlike large models, small models struggled with numerical extraction and calculation. To enhance their reasoning capabilities, we developed:
  • CoT-Supervised Fine-Tuning (CoT-SFT): A three-step process where the model generates relevant formulas, extracts parameter values and units from questions, and uses Numbat for final calculation.
  • Data Augmentation (DA): Using Qwen-max to generate and verify new training examples with correct reasoning steps. Details are provided in Appendix A.2.3.
@@ 4.3 Evaluation Metrics
We utilized Numbat to evaluate the prediction generated by each model against the gold-standard answer. A prediction was deemed correct if the relative error, (prediction - gold) / gold, was less than 1%. We used accuracy, which is the proportion of questions correctly answered, as our metric.
@@ 5 Experimental Results
We present the evaluation results on the Chinese version of FormulaReasoning. The results on the English version are in Appendix A.3.1, which are consistent with those on the Chinese version. The results of PoT are in A.3.2, not better than CoT.
@@
Model HoF test HeF test Avg. GLM-4.5-Air 59.32 61.50 60.37 ERNIE-4.5-21B-A3B 66.34 67.96 67.12 Gemma-3n-E4B-it 14.04 14.47 14.25 Llama-3.3-8b-Instruct 8.47 9.56 9.00 Mistral-Small-3.2-24b-Instruct 74.33 72.35 73.37 GPT-oss-20B 85.96 82.95 84.50 Gemini-2.5-Flash-Lite 81.11 81.40 81.25 Seed-OSS-36B-Instruct 81.84 84.24 83.00 GPT-4o 69.49 65.37 67.50 GLM-4-plus 61.74 63.57 62.63 Qwen3-235B-A22B-Thinking-2507 80.63 83.98 82.25 DeepSeek-R1 88.62 91.47 90.00 o3-mini 92.01 89.66 90.87 GPT-5 92.01 92.76 92.37 Human 93.49 90.47 92.03
@@ 5.1 Results of LLMs with CoT Prompts
We evaluated a diverse range of LLMs across different parameter scales on FormulaReasoning. Our analysis reveals that model performance correlates with both parameter scale and architectural design. As shown in Table 4, large-scale models (e.g., DeepSeek-R1 and o3-mini) achieved competitive results with accuracies exceeding 88% on both HoF and HeF metrics. However, substantial performance gaps existed for smaller-scale models, where architectures like Llama-3.3-8B and Gemma-3N-E4B-it exhibit accuracies below 15%, while even mid-sized models like GLM-4-Air only reached around 60%. Notably, despite their impressive capabilities, even advanced large-scale models such as GPT-4o (67.50%), GLM-4-plus (62.63%), and Qwen3-235B (82.25%) still trailed behind human performance (92.03%). This significant variance suggests that while parameter scaling yields notable improvements, there remains considerable optimization potential for resource-efficient models. Future work should investigate architectural adaptations and training strategies specifically tailored for compact models, which are critical for practical deployment scenarios where computational resources are constrained. The observed performance spectrum underscores the importance of scale-aware evaluation in formula-based reasoning research.
We provide an error analysis in Appendix A.4.
PRM scores to assess the reasoning steps generated by LLMs are included in Appendix A.3.3.
@@ 5.2 Results of LLMs with Formula Retriever
We found that incorporating retrieved formulas into the prompts improved the performance of LLMs. For GLM-4-Air, the accuracy increased by 1.94 from 65.13 to 67.07 on the HoF. Note that it would be meaningless to carry out this experiment on the HeF test set where the required formulas were not included when training the retriever.
Our investigation found that the top five formulas retrieved often included irrelevant ones, as the number of formulas required varied for different questions. The presence of extraneous formulas might introduce noise, suggesting considerable room for improvement in structured formula retrieval.
@@ 5.3 Results of Supervised Fine-Tuned Models
The results of supervised fine-tuned small models are included in Appendix A.3.4. Both CoT-SFT and DA helped improve accuracy. In particular, Qwen2.5-Math-7B-Instruct with CoT-SFT outperformed GPT-4o, demonstrating the effectiveness of our fine-tuning method: delegating numerical calculations to a calculator and focusing on CoT reasoning.
@@ 5.4 Results of DPO (including Results on Other Numerical Reasoning Datasets)
As shown in Appendix A.3.5, DPO with the preference data derived from the training set of FormulaReasoning generally improved the accuracy of 1.5B–7B models on both FormulaReasoning and other numerical reasoning datasets: GSM8K (Cobbe et al., 2021a), MATH Hendrycks et al. (2021), SVAMP Patel et al. (2021), and GaoKao2023-en Chen et al. (2025). These models exhibited much lower accuracy on FormulaReasoning than those datasets. The results underscore the distinct characteristics of our dataset compared to other numerical reasoning benchmarks and its additional value as a potential DPO resource.
6 Conclusion
We introduced FormulaReasoning, a bilingual dataset for formula-based numerical reasoning. For each question, we annotated its complex reasoning steps with normalized physics formulas. We constructed a consolidated formula database after merging equivalent formulas, serving as an external knowledge base to be used in RAG. We used our dataset to evaluate LLMs of various sizes, RAG, fine-tuned small models, and DPO methods, revealing significant performance variations across model scales. Our findings highlight that while larger models show promising performance, smaller-scale LLMs still face significant challenges in formula-based numerical reasoning, indicating opportunities for further advancements in multi-step reasoning guided by domain knowledge.
@@ 6 Conclusion
Limitations
One limitation of our work is that although FormulaReasoning provides step-level supervision information (i.e., formulas), due to the diversity of reasoning paths and formula expressions, we only used PRM to perform step-level process evaluation. Another limitation is that our dataset is focused on physics. We chose junior high school physics because it is not too hard to be understood by ordinary people, which benefited our annotation and evaluation efforts. It is possible to explore formula-based numerical reasoning in other domains such as chemistry and engineering.
@@ A.2.2 Formula Retriever
Each question might correspond to multiple correct formulas, and we ensured that the same question did not appear twice in the same batch when loading the data. Based on the implementation of Chinese-BERT-wwm-base, we tested the retrieval performance on the HoF test set and found that Recall@5 reached 82.46%.
@@ A.2.5 Implementation Details
We accessed DeepSeek-R1 through the DeepSeek API [^5] https://api-docs.deepseek.com/, Qwen series models through the Alibaba Bailian platform [^6] https://bailian.console.aliyun.com/, and other models through OpenRouter [^7] https://openrouter.ai/ with their default hyper-parameters. Qwen2.5-Math-7B/1.5B-Instruct were fine-tuned on NVIDIA GeForce RTX 3090 GPUs with Huggingface Transformers and Pytorch 2.3, where we used a learning rate of 1e-4 and a batch size of 8 to fine-tune them. For Qwen2.5-Math-7B-Instruct model, we fine-tune it with LoRA (Hu et al., 2021) with r=8, alpha=32. For Qwen2.5-Math-1.5B-Instruct model, we fine-tune it with full fine-tuning. We tested the models after training for 3 epochs which is enough for training. We set temperature=0 for direct answer generation and set the maximum output length to 1,024. In the case of CoT-SFT, which directly outputted formulas along with numerical values and units of parameters, we set temperature=0 to obtain the response with the highest probability and set the maximum output length to 512 to obtain formulas and parameters. Training Qwen2.5-Math-7B-Instruct(lora) and Qwen2.5-Math-1.5B-Instruct(full) in CoT-SFT used 3 and 8 hours, respectively.
@@ A.3.1 English Version of FormulaReasoning
The evaluation results for LLMs with CoT prompts on the English FormulaReasoning dataset are presented in Table 5. Large-scale models (e.g., GPT-5 and DeepSeek-R1) achieved competitive results with accuracies exceeding 87% on both HoF and HeF metrics. However, substantial performance gaps existed for smaller-scale models, where architectures like Gemma-3n-E4B-it and Llama-3.3-8b-Instruct exhibited accuracies below 35%, while even mid-sized models like GLM-4.5-Air only reached around 73%. We found that Gemma-3n and Qwen3-235B showed performance drops compared to their results on the Chinese version due to instruction-following issues in the English version, where these models failed to output required units as specified in the prompt.
Notably, despite their impressive capabilities, even the best model GPT-5 (88.63%) still trailed behind human performance (92.03% as reported in the main text), falling short by 3.4 percentage points. Other advanced models such as GPT-4o (75.38%) and GLM-4-plus (72.37%) showed even larger gaps with human performance.
This significant variance suggests that while parameter scaling yields notable improvements, there remains considerable optimization potential for resource-efficient models. Future work should investigate architectural adaptations and training strategies specifically tailored for compact models, which are critical for practical deployment scenarios where computational resources are constrained. The observed performance spectrum underscores the importance of scale-aware evaluation in formula-based reasoning research.
@@
Model HoF test HeF test Avg. GLM-4.5-Air 72.64 74.42 73.53 ERNIE-4.5-21B-A3B 63.20 56.85 60.13 Gemma-3n-E4B-it 2.66 2.84 2.75 Llama-3.3-8b-Instruct 32.45 31.52 32.00 Mistral-Small-3.2-24b-Instruct 70.70 68.99 69.87 GPT-oss-20B 86.20 84.75 85.48 Gemini-2.5-Flash-Lite 75.30 76.49 75.88 Seed-OSS-36B-Instruct 75.30 71.58 73.50 GPT-4o 75.79 74.94 75.38 GLM-4-plus 75.79 68.73 72.37 Qwen3-235B-A22B-Thinking-2507 62.47 66.67 64.50 DeepSeek-R1 87.17 87.60 87.38 o3-mini 84.02 82.17 83.13 GPT-5 88.38 88.89 88.63
A.3.2 PoT Prompts
@@
GPT-4o (CoT) 69.49 65.37 67.50
GPT-4o (PoT) 54.72 41.34 48.25
A.3.3 PRM Scores
Tables 7 and 8 show the PRM scores in one-step and multi-step settings, respectively.
Table 7: PRM score (one-step) of LLMs with CoT prompts.
Model HoF test HeF test Avg. GLM-4.5-Air 0.9301 0.9436 0.9369 ERNIE-4.5-21B-A3B 0.7753 0.7582 0.7668 Gemma-3n-E4B-it 0.3695 0.3660 0.3678 Llama-3.3-8b-Instruct 0.2984 0.2408 0.2696 Mistral-Small-3.2-24b-Instruct 0.8576 0.8133 0.8355 GPT-oss-20B 0.5704 0.5203 0.5454 Gemini-2.5-Flash-Lite 0.5821 0.5818 0.5820 Seed-OSS-36B-Instruct 0.8766 0.8764 0.8765 GPT-4o 0.6150 0.5583 0.5867 GLM-4-plus 0.6728 0.6591 0.6660 Qwen3-235B-A22B-Thinking-2507 0.6502 0.6476 0.6489 DeepSeek-R1 0.7395 0.7393 0.7394 o3-mini 0.7537 0.7556 0.7547 GPT-5 0.8196 0.8249 0.8223
Table 8: PRM score (multi-step) of LLMs with CoT prompts.
Model HoF test HeF test Avg. GLM-4.5-Air 0.9336 0.9438 0.9387 ERNIE-4.5-21B-A3B 0.8695 0.8711 0.8703 Gemma-3n-E4B-it 0.7366 0.7360 0.7363 Llama-3.3-8b-Instruct 0.5507 0.5288 0.5398 Mistral-Small-3.2-24b-Instruct 0.9427 0.9438 0.9433 GPT-oss-20B 0.5738 0.5266 0.5502 Gemini-2.5-Flash-Lite 0.8621 0.8719 0.8670 Seed-OSS-36B-Instruct 0.8789 0.8779 0.8784 GPT-4o 0.8655 0.8444 0.8550 GLM-4-plus 0.8726 0.8738 0.8732 Qwen3-235B-A22B-Thinking-2507 0.7946 0.7996 0.7971 DeepSeek-R1 0.7603 0.7535 0.7569 o3-mini 0.8104 0.8032 0.8068 GPT-5 0.8196 0.8248 0.8222
Table 9: Accuracy of supervised fine-tuned models.
Model HoF test HeF test Avg.
Qwen2.5-Math-7B-Instruct 65.62 58.91 62.27
   + CoT-SFT 72.64 +7.02+7.02 65.63 +6.72+6.72 69.14 +6.87+6.87
   + DA 67.55 +1.93+1.93 62.53 +3.62+3.62 65.04 +2.77+2.77
Qwen2.5-Math-1.5B-Instruct 64.41 55.56 59.99
   + CoT-SFT 66.34 +1.93+1.93 53.75 1.81-1.81 60.05 +0.06+0.06
   + DA 66.59 +2.18+2.18 59.17 +3.61+3.61 62.88 +2.89+2.89